Noluvpablo:

What is the exact distance from (−5, 1) to (3, 0)?

WendyMorales2007:

d = √(x2 - x1)² + (y2 - y1)² given two points, (-5, 1) (3, 0) d = √(3 - (-5)² + (0 - 1)² d = √(8)² + (-1)² d = √64 + 1 d = √65

HelpASAP:

@WendyMorales2007 yeh thats what i got. I wasnt sure if i was meant to type out the work tho lol